Circular No. 7769

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 51P/HARRINGTON P. Manteca, Observatorio de Begues (near Barcelona), reports that his

observations on Dec. 6.1 UT (0.31-m f/6.3 Schmidt-Cassegrain reflector + CCD) showed this comet to have split, the two components being separated by some 10′′ on an east-west line. Each around mag 17.0–17.4, the western component was perhaps up to 0.4 mag brighter than the eastern, which is evidently the one that was under observation during July-November. On Dec. 7.0 and 7.9 Manteca gave m_{1} = 16.4 for the eastern component and m_{1} = 16.6 for the western. The eastern component is also clearly the object defined as component A at the comet’s 1994 apparition, when two much fainter components, B and C, were also recorded with an effective perihelion time almost 0.3 day later (cf. IAUC 6089). The current western component is therefore to be denoted as component D, and its effective perhelion time on 2001 June 5 is only ∼ 0.006 day later than that of component A. The splitting was confirmed by R. Naves and M. Campas (Observatorio Montcabre, also near Barcelona; 0.25-m f/3.3 Schmidt-Cassegrain reflector + CCD), who gave m_{1} as 16.9 for component A and 16.7 for component D on Dec. 6.9 and 16.2 for component A and 16.3 for component D on Dec. 7.9. R. Ferrando (Observatorio Pla D’Arguines, near Valencia; 0.30-m Schmidt-Cassegrain reflector + CCD) gave m_{1} as 16.5 for component A and 16.7 for component D on Dec. 7.9. The comet is noticeably brighter than the predicted magnitude given in the ephemeris on MPC 43692.

4U 0142+61 B. Kern and C. Martin, California Institute of Technology, communi-

cate: “We observed the anomalous x-ray pulsar (AXP) 4U 0142+61 on Nov. 16–17 UT with the Palomar 5-m telescope, using an unfiltered CCD observing in a periodic frame-transfer mode, synchronized with the known 8.7-s x-ray period (Gavriil and Kaspi 2001, http://xxx.lanl.gov/abs/astroph/?0107422). In 10-phase bins, we find optical pulsations with a 30percent pulsed fraction. The pulse profile is similar in morphology and agrees within 0.1 in phase with the soft x-ray pulse profile. This is the first detection of optical pulsations in an AXP. Our discovery unambiguously confirms the optical counterpart (Hulleman et al. 2000, Nature 408, 689), and the large pulsed fraction rules out x-ray reprocessing as the source of pulsed optical emission.”

2001 December 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green